# Detecting environmental effects in gravitational waves from binaries perturbed by periodic forces

### v2

@@ I Introduction
Astrophysical sources of gravitational waves (GW) are likely to form in high density environments, such as star clusters, nuclear clusters, and accretion or circumbinary discs. Understanding and modelling the physical mechanisms at play in these environments is crucial, as they are responsible for determining the efficiency of the various formation channels for black hole binaries (BHB), and ultimately result in the expected event rates of detectable signals for current and future GW detectors [see the following reviews and white papers, among the innumerable relevant works, 1, 46, 85, 4, 47, 33]. In addition to determining the rates, different astrophysical environments set characteristic ranges and correlations in the expected parameters of observed BHBs, as exemplified by the numerous efforts to link LIGO/Virgo/Kagra observations to specific formation channels [16, 91, 110, 6, 142, 71, 154, 73, 117].
@@
In almost all astrophysically motivated scenarios, the direct influence of the source’s environment on the emitted GWs is expected to be extremely small once BHBs reach the separations required to produce detectable GWs [with the possible exception of Pulsar Timing Array sources, see, e.g., 65, 119]. However, extremely small need not necessarily mean negligible. These influences are referred to with the umbrella term of ”environmental effects” (EE), which are defined as small perturbations with respect to the expected vacuum GWs that arise as a consequence of the source’s astrophysical environment. Common examples of EEs include dynamical perturbations caused by third bodies, effects of peculiar velocity and other accelerations, as well as the study of frictional forces motivated by both gaseous and dark matter backgrounds [13, 68, 88, 20, 135, 99, 49, 80, 149, 62, 32, 28, 155, 131].
@@
Since the early mentions in the nineties [25, 108] to the seminal work of the mid two-thousands [78, 74, 12], the study of EE, i.e. astrophysically motivated perturbations to GW emission, has been gaining traction due to the ever closer dawn of space-based GW detectors sensitive to the milli-Hz band [in particular the recently adopted LISA mission, 33]. The scientific motivations are threefold: Firstly, the detection of even a single source that is clearly perturbed by EEs would allow to set unprecedented constraints on the surroundings of BHBs and the physical processes that facilitate their formation and shape their evolution [41, 38, 156, 124, 32, 42, 62, 114]. Secondly, the presence of unmodelled EEs may produce biases in the recovered binary parameters (see references above). Thirdly, deviations from the expected vacuum evolution of sources may be mislabelled as resulting from non-standard physics, in the case that astrophysical influences are not modelled correctly [see, e.g., 90, 151, 23, 22, 32, 155]. The latter point is especially relevant when interpreting signals from extreme mass-ratio inspirals, which are supposed to provide extremely precise measurement of the metric in the close vicinity of massive BHs [11, 60, 3, 97]. In short, the careful study of EE is is required to both preserve and broaden the scientific goals of future GW detectors in the joint fields of fundamental physics and astrophysics.
@@
where SNR is the signal to noise ratio of the given GW, can significantly reduce the match between a waveform template and the data and reduce the effectiveness of parameter estimation [see, e.g., 34, 35, 74, 70, 92]. Secondly, most environmental effects may indeed be modelled with simple secular de-phasing prescriptions, at least as a first approximation [see 155, for an impression on the broad relevance of dephasing in the context of EE]. As an example, any coupling between the environment’s and the binary’s energy will cause a de-phasing by the following relation:
@@
where $\Delta\dot{f}$ is a change in the binary’s chirp due to the coupling. Additionally, any kind of EE that involves a Doppler shift [see, e.g., 26, where modifications to the GW phase and GW amplitude are discussed in detail], such as peculiar motion or gravitational red-shift, is also well approximated via a dephasing [88, 68, 135, 29, 152]:
@@
Secular de-phasing is often considered the most promising avenue to detect EE in GWs, and several works have used the reference value of Eq. 1 to claim detectability [^1] Including, admittedly, the corresponding author’s own work.. While the latter criterion can provide a good estimate in certain limits, several careful parameter inference studies have highlighted how de-phasing suffers from many-fold intrinsic limitations which may greatly hinder the prospective of correctly identifying it in realistic binary sources [24, 72, 92]. Firstly, secular dephasing is preferentially accumulated in the very early inspiral stages of a binary’s observable evolution, where the binary only chirps slowly. This corresponds to a rather low available signal-to-noise (SNR) in which the perturbation may effectively be constrained. Secondly, tracking an accumulated phase shift over a very long inspiral may bee exceedingly hard due to the intrinsic challenges of matched-filtering [53]. Thirdly, secular de-phasing is similar to many other post-Newtonian (PN) corrections to the binary’s vacuum evolution, which also add slowly accumulating contributions to the total GW phase [see e.g. 18]. Finally, and perhaps most crucially, it is the case that a hypothetical constant dephasing term, i.e. a phase offset, has exactly the same form as the source’s initial phase (as can be seen at a glance in, e.g., Eq. 28), leading to potential degeneracies. The consequence of this is that only the rate of variation in the de-phasing is truly informative, rather than the magnitude of the de-phasing itself.
@@ II.1 Motion of a circular binary under time varying forces
We consider the case of a binary source of GWs with arbitrary total mass $M$ and reduced mass $\mu$. The binary’s orbit is described by the solutions to the six osculating elements $\left[p,e,\nu,\omega,\Omega,I\right]$ [^2] Semi-latus rectum, eccentricity, true anomaly, argument of peri-apsis, longitude of ascending node and inclination., which fix the coordinate position of the reduced mass in space [following here the formalism detailed in 96]:
@@
In this work, we restrict our analysis to in-plane perturbative forces, as out of plane displacements will only ever affect the orbital inclination angles. Note that, if sufficiently strong, variations in the inclination angles may well produce detectable modulations in the relative amplitude of the GW’s plus and cross polarisations. We leave this effect for future investigation. We treat the binary’s orbit in the Newtonian limit. While this assumption is known to break down as the orbital velocity approaches relativistic values, it is justifiable by two arguments: First, most EEs will affect the binary more strongly at large separations, where orbital speeds are low and GW emission is weaker. Second, the assumption is justified in a strictly perturbative sense, as additional cross terms [^3] Cross terms between environmental perturbations and post Newtonian dynamics have been extensively studied in [145, 146], and have interesting consequences on the secular evolution of triple systems and orbits in accretion discs [157]. between the environmental perturbations and relativistic dynamics will be suppressed with respect to the leading order terms in $1/c^{2}$. Note that under these assumptions, the binary’s motion will be confined to a unique plane, which we identify as the $(x,y)$ plane.
@@
where the complex coefficients $B^{\rm T}_{n}$ and $B^{\rm R}_{n}$ are only functions of the binary’s separation and their magnitude and dependence on $r$ encodes the forces that are applied on the binary by the given EE. The form of Eqs. II.1 is to be understood as a Fourier series in the limit in which the binary residency time at the separation $r$ is much larger than one orbital period, i.e. the adiabatic limit. In that case, we can also allow (with a slight abuse of notation) the parameter $n$ to take non-integer values in order to describe periodic perturbations that do not exactly match a multiple of the system’s dynamical time. We can insert the perturbing force into the Lagrange planetary equations to find analytical solutions to the orbital elements. As it will be directly related to GW observables, we report the explicit first order perturbative solution [^4] Here we take the chance to note an imprecision present in [156], in which the computations did not properly propagate the perturbation in the Kepler equation and resulted in a mistake in Eq. (33). However, the latter does not change the qualitative results in [156], as the physical scaling remains unaffected. of the true longitude $\theta=\nu+\omega$:
@@
$\delta\theta_{n}^{\rm R}$ $=-\frac{B_{n}^{\rm R}p_{0}^{2}\exp(i2\pi nf_{\rm K}t)}{GM\left(n^{6}-5n^{4}+4n^{2}\right)}n\Big{(}i(-4+n^{2})4$
@@
where the perturbations (denoted by $\delta$) are purely oscillatory and indeed do not cause any secular evolution of the binary over timescales longer than the forcing. We also report the explicit solution for the binary’s separation, which oscillates about its unperturbed value:
@@
We find, as expected, that oscillatory forces cause a corresponding behaviour in the orbital elements with a periodicity comparable to the forcing. A visualisation of the resulting binary motion can be seen in Figure 1. In essence, Eqs. 12 and II.1 state that a periodic, in-plane perturbation with an amplitude $B$ and with a periodicity of $nf_{\rm K}$ will produce variations in the orbital phase of typical order $Bp_{0}^{2}/(GM)$ at the different harmonics $(2-n,n,2+n)$. The amplitude of high-frequency fluctuations is suppressed as $1/n^{2}$, due to the shorter available time to apply an acceleration. We can summarise the results schematically as follows:
Force $\to$ Phase $\big{\{}B,[n]f_{\rm K}\big{\}}$to $\Big{\{}\sim\frac{Bp_{0}^{2}}{GM},[2-n,n,2+n]f_{\rm K}\Big{\}}$
In Fig. 2 we show the dependence of the true longitude perturbation coefficients on the periodicity $n$. We find that, when neglecting some special values of $n$ as discussed below, the coefficients for tangential and radial perturbations can be well fit by the simpler expressions $3/n^{2}$ and $1/(n+n^{2})$, respectively.
@@ II.2 Vacuum and perturbed waveform models
$\psi$ $=-\pi/4+\phi_{\rm c}-2\pi ft+2\left(8\pi\mathcal{M}f\right)^{-5/3}$ (28)
@@
Eqs. 27 and 28 represent a post-Newtonian, quasi-circular waveform model that can describe the effect of arbitrary time-varying forces on the binary’s phase. The number of additional free parameters with respect to a vacuum waveform is limited to the number of free parameters required to specify the perturbations on the binary. For clarity’s sake, let us summarise again the assumptions that went into the construction of the waveform model:
@@
Given these assumptions, we propose our perturbed waveform model as a proxy for more sophisticated templates that can account for both strong gravity effects and environmental perturbations in a fully relativistic way. We highlight the exciting prospect of measuring forces directly from, e.g., hydrodynamical and GRMHD simulations of gas-embedded binaries or precise integrations of few body dynamical scatterings, including them in Eqs. 27 and 28, and performing parameter estimation directly on such numerical results.
@@ III SNR criteria
where the prefactor 8 arises from the normalisation of the one sided power spectrum [101]. Note that the latter equation is often only a rough estimate of the detectability of a waveform perturbation. In particular, the commonly chosen criterion $\delta$SNR ¿8 does not account for degeneracies and correlations between the binary vacuum parameters and the perturbations under scrutiny. Thus, Eq. 35 often overestimates the detectability of secular de-phasing, which is strongly degenerate with both the binary’s initial phase as well as other parameters such as effective spin [^6] This particular degeneracy is in fact responsible for the mismatch between the Bayesian analysis performed in this work and recent works [see, e.g., 61] on gas de-phasing based on the pycbc package, as the latter treats the initial phase as an extrinsic parameter., as seen in Fig. 7. For this reason, we will test the simple detectability criteria with several MCMC parameter recovery tests. From here on, we will refer to the results of Eq. 35 as a ”$\delta$SNR” estimate.
@@
In this paper we perform several numerical parameter recovery tests to validate the simpler analytical prescriptions of the previous sections. We adopt a vacuum waveform model of the type described by Eqs. 27 and 28. We include the phase evolution of a non-spinning binary up to third PN order by using the formulas from [19]. Similarly, we also account for all higher order modes beyond quadrupolar GW emission up to order $x^{5/2}$, where $x$ is the post-Newtonian parameter. Following [34], we also include the leading order effect of spin through the effective parameter $\mathcal{B}$, which enters the phase at first-and-a-half post-Newtonian order and represents a projection of the orbital angular momentum and the individual binary component spins. In this work, we do not include any spin evolution term nor the effects of spin-spin coupling. Finally, we include the effect of periodic dephasing by means of the baseline model detailed above, for a single periodicity. While this suffices to obtain correct detectability estimates, we stress that the latter cannot replace the full formulas for true longitude perturbations in actual applications.
@@ IV.1 Detectability for monochromatic sources
where we used the definition of the signal-to-noise ratio for a monochromatic unperturbed source [see, e.g., 19, 83, 101]. We can also easily extrapolate the results of performing the integral in Eq. 50 for a full spectrum of perturbations. Re-introducing the subscript $n$ to denote the various periodic force components, we find that the total SNR grows in quadrature:

### v3

# A novel category of environmental effect in gravitational waves from binaries perturbed by periodic forces

@@ I Introduction
Astrophysical sources of gravitational waves (GW) are likely to form in high density environments, such as star clusters, nuclear clusters, and accretion or circumbinary discs. Understanding and modelling the physical mechanisms at play in these environments is crucial, as they are responsible for determining the efficiency of the various formation channels for black hole binaries (BHB), and ultimately result in the expected event rates of detectable signals for current and future GW detectors [see the following reviews and white papers, among the innumerable relevant works, 1, 49, 91, 4, 50, 34]. In addition to determining the rates, different astrophysical environments set characteristic ranges and correlations in the expected parameters of observed BHBs, as exemplified by the numerous efforts to link LIGO/Virgo/Kagra observations to specific formation channels [16, 97, 116, 6, 148, 75, 161, 77, 123].
@@
In almost all astrophysically motivated scenarios, the direct influence of the source’s environment on the emitted GWs is expected to be extremely small once BHBs reach the separations required to produce detectable GWs [with the possible exception of Pulsar Timing Array sources, see, e.g., 69, 125]. However, extremely small need not necessarily mean negligible. These influences are referred to with the umbrella term of ”environmental effects” (EE), which are defined as small perturbations with respect to the expected vacuum GWs that arise as a consequence of the source’s astrophysical environment. Common examples of EEs include dynamical perturbations caused by third bodies, effects of peculiar velocity and other accelerations, as well as the study of frictional forces motivated by both gaseous and dark matter backgrounds [13, 72, 94, 20, 141, 106, 52, 86, 156, 66, 33, 29, 162, 137, 56].
@@
Since the early mentions in the nineties [26, 114] to the seminal work of the mid two-thousands [83, 78, 12], the study of EE, i.e. astrophysically motivated perturbations to GW emission, has been gaining traction due to the ever closer dawn of space-based GW detectors sensitive to the milli-Hz band [in particular the recently adopted LISA mission, 34]. The scientific motivations are threefold: Firstly, the detection of even a single source that is clearly perturbed by EEs would allow to set unprecedented constraints on the surroundings of BHBs and the physical processes that facilitate their formation and shape their evolution [44, 41, 163, 130, 33, 45, 66, 120]. Secondly, the presence of unmodelled EEs may produce biases in the recovered binary parameters (see references above). Thirdly, deviations from the expected vacuum evolution of sources may be mislabelled as resulting from non-standard physics, in the case that astrophysical influences are not modelled correctly [see, e.g., 96, 158, 24, 23, 33, 162]. The latter point is especially relevant when interpreting signals from extreme mass-ratio inspirals, which are supposed to provide extremely precise measurement of the metric in the close vicinity of massive BHs [11, 64, 3, 103]. In short, the careful study of EE is required to both preserve and broaden the scientific goals of future GW detectors in the joint fields of fundamental physics and astrophysics.
@@
where SNR is the signal to noise ratio of the given GW, can significantly reduce the match between a waveform template and the data and reduce the effectiveness of parameter estimation [see, e.g., 35, 36, 78, 74, 98]. Secondly, most environmental effects may indeed be modelled with simple secular de-phasing prescriptions, at least as a first approximation [see 162, for an impression on the broad relevance of dephasing in the context of EE]. As an example, any coupling between the environment’s and the binary’s energy will cause a de-phasing by the following relation:
@@
where $\Delta\dot{f}$ is a change in the binary’s chirp due to the coupling. Additionally, any kind of EE that involves a Doppler shift [see, e.g., 27, where modifications to the GW phase and GW amplitude are discussed in detail], such as peculiar motion or gravitational red-shift, is also well approximated via a dephasing [94, 72, 141, 30, 159]:
@@
Secular de-phasing is often considered the most promising avenue to detect EE in GWs, and several works have used the reference value of Eq. 1 to claim detectability [^1] Including, admittedly, the corresponding author’s own work.. While the latter criterion can provide a good estimate in certain limits, several careful parameter inference studies have highlighted how de-phasing suffers from many-fold intrinsic limitations which may greatly hinder the prospective of correctly identifying it in realistic binary sources [25, 76, 98]. Firstly, secular dephasing is preferentially accumulated in the very early inspiral stages of a binary’s observable evolution, where the binary only chirps slowly. This corresponds to a rather low available signal-to-noise (SNR) in which the perturbation may effectively be constrained. Secondly, tracking an accumulated phase shift over a very long inspiral may bee exceedingly hard due to the intrinsic challenges of matched-filtering [57]. Thirdly, secular de-phasing is similar to many other post-Newtonian (PN) corrections to the binary’s vacuum evolution, which also add slowly accumulating contributions to the total GW phase [see e.g. 18]. Finally, and perhaps most crucially, it is the case that a hypothetical constant dephasing term, i.e. a phase offset, has exactly the same form as the source’s initial phase (as can be seen at a glance in, e.g., Eq. II.2), leading to potential degeneracies. The consequence of this is that only the rate of variation in the de-phasing is truly informative, rather than the magnitude of the de-phasing itself.
@@ II.1 Motion of a circular binary under time varying forces
We consider the case of a binary source of GWs with arbitrary total mass $M$ and reduced mass $\mu$. The binary’s orbit is described by the solutions to the six osculating elements $\left[p,e,\nu,\omega,\Omega,I\right]$ [^2] Semi-latus rectum, eccentricity, true anomaly, argument of peri-apsis, longitude of ascending node and inclination., which fix the coordinate position of the reduced mass in space [following here the formalism detailed in 102]:
@@
In this work, we restrict our analysis to in-plane perturbative forces, as out of plane displacements will only ever affect the orbital inclination angles. Note that, if sufficiently strong, variations in the inclination angles may well produce detectable modulations in the relative amplitude of the GW’s plus and cross polarisations. We leave this effect for future investigation. We treat the binary’s orbit in the Newtonian limit. While this assumption is known to break down as the orbital velocity approaches relativistic values, it is justifiable by two arguments: First, most EEs will affect the binary more strongly at large separations, where orbital speeds are low and GW emission is weaker. Second, the assumption is justified in a strictly perturbative sense, as additional cross terms [^3] Cross terms between environmental perturbations and post Newtonian dynamics have been extensively studied in [153, 152], and have interesting consequences on the secular evolution of triple systems and orbits in accretion discs [164]. between the environmental perturbations and relativistic dynamics will be suppressed with respect to the leading order terms in $1/c^{2}$. However, see [22] for a fully relativistic approach appropriate for high mass-ratio binaries. Note that under these assumptions, the binary’s motion will be confined to a unique plane, which we identify as the $(x,y)$ plane.
@@
where the complex coefficients $B^{\rm T}_{n}$ and $B^{\rm R}_{n}$ are only functions of the binary’s separation and their magnitude and dependence on $r$ encodes the forces that are applied on the binary by the given EE. The form of Eqs. II.1 is to be understood as a Fourier series in the limit in which the binary residency time at the separation $r$ is much larger than one orbital period, i.e. the adiabatic limit. In that case, we can also allow (with a slight abuse of notation) the parameter $n$ to take non-integer values in order to describe periodic perturbations that do not exactly match a multiple of the system’s dynamical time. We can insert the perturbing force into the Lagrange planetary equations to find analytical solutions to the orbital elements. As it will be directly related to GW observables, we report the explicit first-order perturbative solution [^4] Here we take the chance to note an imprecision present in [163], in which the computations did not properly propagate the perturbation in the Kepler equation and resulted in a mistake in Eq. (33). However, the latter does not change the qualitative results in [163], as the physical scaling remains unaffected. of the true longitude $\theta=\nu+\omega$:
@@
$\delta\theta_{n}^{\rm R}$ $=-\frac{B_{n}^{\rm R}p_{0}^{2}\exp(i2\pi nf_{\rm K}t)}{GM\left(n^{6}-5n^{4}+4n^{2}\right)}n\Big{(}i(-4+n^{2})$
@@
where the perturbations (denoted by $\delta$) are purely oscillatory and vanish when averaged over the true anomaly on timescales longer than the forcing (again for $n\neq[0,1,2]$). We also report the explicit solution for the binary’s separation, which oscillates about its unperturbed value:
@@
We find, as expected, that oscillatory forces cause a corresponding behaviour in the orbital elements with a periodicity comparable to the forcing. A visualisation of the resulting binary motion can be seen in Figure 1. In essence, Eqs. 12 and II.1 state that a periodic, in-plane perturbation with an amplitude $B$ and with a periodicity of $nf_{\rm K}$ will produce variations in the orbital phase of typical order $Bp_{0}^{2}/(GM)$ at the different harmonics $(n-2,n,n+2)$. The amplitude of high-frequency fluctuations is suppressed as $1/n^{2}$, due to the shorter available time to apply an acceleration. We can summarise the results schematically as follows:
Force $\to$ Phase
$\big{\{}B,[n]f_{\rm K}\big{\}}$to $\Big{\{}\sim\frac{Bp_{0}^{2}}{GM},[n-2,n,n+2]f_{\rm K}\Big{\}}$ ,
where the central perturbation corresponds to the direct effect of tangential or radial forces on the binary phase, while the two other arise as a response of the binary’s elements to the changing instantaneous velocity or radial separation, respectively. In Fig. 2 we show the dependence of the true longitude perturbation coefficients on the periodicity $n$. We find that, when neglecting some special values of $n$ as discussed below, the coefficients for tangential and radial perturbations can be well fit by the simpler expressions $3/n^{2}$ and $1/(n+n^{2})$, respectively. Note that these approximations are valid both in the limit $n\to 0$ and $n\to\infty$, and will be used in the numerical tests of section IV.
@@ II.2 Vacuum and perturbed waveform models
$\psi$ $=-\pi/4+\phi_{\rm c}-2\pi ft+2\left(8\pi\mathcal{M}f\right)^{-5/3}$
@@
Eqs. 27 and II.2 represent a post-Newtonian, quasi-circular waveform model that can describe the effect of arbitrary time-varying forces on the binary’s phase. The number of additional free parameters with respect to a vacuum waveform is limited to the number of free parameters required to specify the perturbations on the binary. For clarity’s sake, let us summarise again the assumptions that went into the construction of the waveform model:
@@
Given these assumptions, we propose our perturbed waveform model as a proxy for more sophisticated templates that can account for both strong gravity effects and environmental perturbations in a fully relativistic way. We highlight the exciting prospect of measuring forces directly from, e.g., hydrodynamical and GRMHD simulations of gas-embedded binaries or precise integrations of few body dynamical scatterings, including them in Eqs. 27 and II.2, and performing parameter estimation directly on such numerical results.
@@ III SNR criteria
where the prefactor 8 arises from the normalisation of the one sided power spectrum [107]. Note that the latter equation is often only a rough estimate of the detectability of a waveform perturbation. In particular, the commonly chosen criterion $\delta$SNR$>$8 does not account for degeneracies and correlations between the binary vacuum parameters and the perturbations under scrutiny. Thus, Eq. 34 often overestimates the detectability of secular de-phasing, which is strongly degenerate with both the binary’s initial phase as well as other parameters such as effective spin [^6] This particular degeneracy is in fact responsible for the mismatch between the Bayesian analysis performed in this work and recent works [see, e.g., 65] on gas de-phasing based on the pycbc package, as the latter treats the initial phase as an extrinsic parameter., as seen in Fig. 7. For this reason, we will test the simple detectability criteria with several MCMC parameter recovery tests. From here on, we will refer to the results of Eq. 34 as a ”$\delta$SNR” estimate.
@@
In this paper we perform several numerical parameter recovery tests to validate the simpler analytical prescriptions of the previous sections. We adopt a vacuum waveform model of the type described by Eqs. 27 and II.2. We include the phase evolution of a non-spinning binary up to third PN order by using the formulas from [19]. Similarly, we also account for all higher order modes beyond quadrupolar GW emission up to order $x^{5/2}$, where $x$ is the post-Newtonian parameter. Following [35], we also include the leading order effect of spin through the effective parameter $\mathcal{B}$, which enters the phase at first-and-a-half post-Newtonian order and represents a projection of the orbital angular momentum and the individual binary component spins. In this work, we do not include any spin evolution term nor the effects of spin-spin coupling. Finally, we include the effect of periodic dephasing by means of the baseline model detailed above, for a single periodicity. While this suffices to obtain correct detectability estimates, we stress that the latter cannot replace the full formulas for true longitude perturbations in actual applications.
@@ IV.1 Detectability for monochromatic sources
where we used the definition of the signal-to-noise ratio for a monochromatic unperturbed source [see, e.g., 19, 89, 107]. We can also easily extrapolate the results of performing the integral in Eq. 49 for a full spectrum of perturbations. Re-introducing the subscript $n$ to denote the various periodic force components, we find that the total SNR grows in quadrature: